OpenStudy (anonymous):

What is the solution to the system of equations graphed below? y = x + 3 y = 3x + 1 A. (1, 4) B. (1, 3) C. (0, 1) D. (0, 3)

Parth (parthkohli):

x + 3 = 3x + 1 x + 2 = 3x 2x = 2 x =1 y = 1 + 3 = 4

Parth (parthkohli):

A. 1,4